If the time constant of the LR circuit illustrated below is τ=5s, what is L, the unknown inductance?

xid-24416354_2

a.

2H

b.

19H

c.

7H

d.

20H

e.

18H
Answer #1

time constant in LR circuit is T = L/R

given t = 5 s

L = T*R

L = 5*R